OpenStudy (anonymous):

When they work together, Stuart and Lucy can deliver flyers to all the hokes in their neighbourhood in 42 min. When Lucy works alone, she can finish the deliveries in 13 min less time than Stuart can when he works alone. When Staurt works alone, how long does he take to deliver the flyers?

hero (hero):

Formula for working together: \[\frac{S \times L}{S + L} = t \] S = Stuart's time L = Lucy's time t = Time it takes when Stuart and Lucy work together In this case: S = x L = x - 13 t = 42 Thus: \[\frac{x(x-13)}{x + x - 13} = 42\] After solving for x, you should get: x = 91

OpenStudy (anonymous):

can u explain the formula to me please?

hero (hero):

This is the only way I can explain it: Michael can complete a task in 5 minutes. Nancy can complete the same task in 7 minutes. How much time would it take if they did the job together? Using the "Time Working Together" formula: \[\frac{ M \times N}{M + N} = t\] Insert the given values M = 5 and N = 7: \[\frac{5 \times 7}{5 + 7} = t\] Now solve for t: \[\frac{35}{12} = t\] \[t = 2.91 \approx 3\] Therefore, if they work together, Michael and Nancy can complete the task in approximately 3 minutes.

hero (hero):

@sabika13

OpenStudy (anonymous):

Thank you so much!!!